OpenStudy (cookiemonster18):

1.For the graph below, what is the approximate y-value of the point of intersection? A.-1 B.4 C.3 D.2

OpenStudy (cookiemonster18):

2.Which of the following systems of inequalities represents the graph? A.y<=2x+4 y>=-x+2 B.2x-y>=4 y<-x+2 C.y>=2x+4 -x+y<=2 D.-2x+y>=4 x+y<2

OpenStudy (cookiemonster18):

For #2 the answer is D im just not sure what the answer to #1 is

OpenStudy (nathalie0918):

ok for #2 this is A

OpenStudy (nathalie0918):

this is B

OpenStudy (nathalie0918):

this is C

OpenStudy (nathalie0918):

this is D

OpenStudy (nathalie0918):

from those graphs you should be able to find out with one fits better for you to match the other graph. Now for #1 do you have a graph?

OpenStudy (cookiemonster18):

okay but whats #1

OpenStudy (nathalie0918):

well do you have the graph for #1?

OpenStudy (cookiemonster18):

For #2 I said the answer was D because when I put it into the graphing calculator it gave me this and I think #1 is D as well but im not sure

OpenStudy (nathalie0918):

well the y-value is closes to 3, so I would go with C

OpenStudy (cookiemonster18):

so 1.C and 2.D?? @nathalie0918

jimthompson5910 (jim_thompson5910):

@CookieMonster18 you wrote this `1.For the graph below, what is the approximate y-value of the point of intersection?` but you forgot to post the graph for problem #1

OpenStudy (cookiemonster18):

im sorry I had to change it because I couldn't post the graph when I made the post so I switched it up a little #1 is really the 2nd and the #2 is really the first question here ill fix it :)

OpenStudy (cookiemonster18):

1.Which of the following systems of inequalities represents the graph? A.y<=2x+4 y>=-x+2 B.2x-y>=4 y<-x+2 C.y>=2x+4 -x+y<=2 D.-2x+y>=4 x+y<2

OpenStudy (cookiemonster18):

2.For the graph above, what is the approximate y-value of the point of intersection? A.-1 B.4 C.3 D.2 @jim_thompson5910

OpenStudy (mathmale):

there appears to be an easily recognized red line in this graph, plus a not-so-easily recognized other line that crosses the red line. Look at where they meet. Using the scale on the x and y axes, estimate the coordinates of the point of intersection. Actually, you need to estimate only the y-coordinate of this point.

OpenStudy (cookiemonster18):

@mathmale can you help me do that

OpenStudy (mathmale):

I see the red line intersecting another line between y=2.0 and y=3.0. Do you?